# NSSI-Net: Multi-Concept Generative Adversarial Network for Non-Suicidal Self-Injury Detection Using High-Dimensional EEG Signals in a Semi-Supervised Learning Framework

### v1

@@ Abstract
Non-suicidal self-injury (NSSI) is a serious threat to the physical and mental health of adolescents, significantly increasing the risk of suicide and attracting widespread public concern. Electroencephalography (EEG), as an objective tool for identifying brain disorders, holds great promise. However, extracting meaningful and reliable features from high-dimensional EEG data, especially by integrating spatiotemporal brain dynamics into informative representations, remains a major challenge. In this study, we introduce an advanced semi-supervised adversarial network, NSSI-Net, to effectively model EEG features related to NSSI. NSSI-Net consists of two key modules: a spatial-temporal feature extraction module and a multi-concept discriminator. In the spatial-temporal feature extraction module, an integrated 2D convolutional neural network (2D-CNN) and a bi-directional Gated Recurrent Unit (BiGRU) are used to capture both spatial and temporal dynamics in EEG data. In the multi-concept discriminator, signal, gender, domain, and disease levels are fully explored to extract meaningful EEG features, considering individual, demographic, disease variations across a diverse population. Based on self-collected NSSI data (n=114), the model’s effectiveness and reliability are demonstrated, with a 7.44% improvement in performance compared to existing machine learning and deep learning methods. This study advances the understanding and early diagnosis of NSSI in adolescents with depression, enabling timely intervention. The source code is available at https://github.com/Vesan-yws/NSSINet.
@@ II-B EEG Studies Exploring NSSI in Depressed Adolescents
In recent years, researchers have increasingly focused on utilizing EEG to explore the neural underpinnings of NSSI. In 2021, a study highlighted significant differences in EEG frequency and spatial components between depressive female adolescents who had attempted suicide and those who engaged in NSSI, demonstrating the potential of EEG as a diagnostic tool to distinguish between these two groups [21]. Moreover, subsequent research has revealed notable alterations in emotion-related EEG components and brain lateralization in response to negative emotional stimuli among adolescents with NSSI, suggesting that these neural changes could be crucial in understanding the emotional dysregulation associated with NSSI [22]. Further investigations have explored the effects of therapeutic interventions on EEG patterns in NSSI patients. For example, Zhao et al. ’s work observed significant changes in EEG microstate patterns before and after repetitive transcranial magnetic stimulation (rTMS) treatment, indicating that EEG can not only reflect the abnormal neural activity related to NSSI but also monitor the impact of therapeutic interventions [23]. These findings underscore the importance of EEG as a non-invasive, real-time measure that can provide insights into the neural dynamics associated with NSSI, potentially leading to more targeted and effective treatments.
@@ III-B Data Acquisition and Preprocessing
A standard EEG preprocessing pipeline was applied to the raw EEG signals to remove artificial artifacts, including eye movements, muscle activity, and environmental noise. Following artifact correction, the sampling rate was downsampled from 500 Hz to 384 Hz, reducing data dimensionality and making the data more efficient for subsequent processing and analysis. To increase the sample size for modeling, each 5-second trial was further segmented into multiple 1-second samples. As a result, the final data format can be represented as 114 participants $\times$ 35 trials $\times$ 5 seconds $\times$ 63 channels $\times$ 384 sampling points.
@@ IV-A Input Data
Considering the data imbalance (NSSI: 77 patients, including 65 females and 12 males; non-NSSI: 37 patients, including 18 females and 19 males) and the gender distribution, a data sampling strategy is implemented across $\mathbb{S}_{l}$, $\mathbb{S}_{u}$, and $\mathbb{T}$. As illustrated in Fig. 2, the sampling process for both the NSSI and non-NSSI groups is explained. For the NSSI group (denoted as DN+), 18 females are randomly selected from the 65 available candidates (represented by the red hatched figures), and all 12 males are included, ensuring equal gender representation within DN+. For the non-NSSI group (denoted as DN-), all 18 females (solid blue figures) are included, along with 12 randomly selected males from the 19 available candidates, achieving a balanced gender distribution comparable to the NSSI group. This sampling process results in a dataset of 30 NSSI patients (DN+) and 30 non-NSSI patients (DN-), balanced across both condition and gender.
@@ IV-B Spatial-Temporal Feature Extraction
To effectively extract spatial and temporal features from input EEG signals, we designed a hybrid neural network that combines CNNs with GRUs. This architecture captures interactions both among different brain regions and across various time points, enhancing the representation of complex EEG dynamics.
@@
To further uncover the relationships of the brain patterns over time and encapsulate the key temporal dynamics of the entire EEG signal sequence, a BiGRU based encoder-decoder structure is designed. The BiGRU-based encoder for temporal information extraction is illustrated in Fig. 2, represented by the green vertical arrow. In our implementation, we used a BiGRU to process the sequence of feature vectors generated by the CNN encoder. The forward and backward passes of the BiGRU produce corresponding hidden states, which are concatenated to form the deep feature representation vector $\mathcal{A}=(a_{1},...,a_{M})\in\mathbb{R}^{M\times 1\times 2F}$. Once $\mathcal{A}$ is obtained from the concatenated hidden states of the BiGRU, it is fed into a Multi-Layer Perceptron (MLP) to derive $\mathcal{Y}=(y_{1},...,y_{M})\in\mathbb{R}^{M\times 1\times F}$. This MLP transforms the feature representation into a more refined feature space, capturing non-linear relationships between the elements of $\mathcal{A}$. Following this, a flatten operation is applied to $\mathcal{Y}$, resulting in the output vector $\mathcal{O}\in\mathbb{R}^{1\times MF}$. The output $\mathcal{O}$ is a rich representation that integrates both spatial features (extracted by the CNN layers) and temporal features (learned through the recurrent layers), providing a comprehensive understanding of the EEG signal’s dynamics.
@@ IV-C2 Gender-Specific Discriminator
Considering the gender bias present in NSSI patients and the variations in EEG pattern representations between females and males, a gender-specific discriminator is designed to capture and leverage gender-specific EEG features. By integrating gender-specific discriminator, the model not only accounts for the gender-based differences but also improves its overall classification accuracy, ensuring that the subtleties in EEG signals attributed to gender are adequately represented and utilized for better prediction outcomes. The loss function of the gender-specific discriminator is given as
@@
Here, $\theta_{G}$ represents the parameters of the feature extractor (generator $G$). $\theta_{e}$ refers to the parameters of the gender classifier $e(\cdot)$. $\mathcal{O}$ refers to the features extracted as described in Section IV-B Spatial-Temporal Feature Extraction, where $\mathcal{O}^{\mathbb{M}}$ and $\mathcal{O}^{\mathbb{F}}$ represent the features of male and female samples, respectively. $N_{\mathbb{M}}$ and $N_{\mathbb{F}}$ denote the number of male and female samples in each batch, respectively. This loss function is designed to ensure that the model effectively captures gender-specific EEG features, which are crucial, particularly in addressing the gender bias present in the data.
@@ IV-C3 Domain-Specific Discriminator
Due to the inherent individual differences in EEG signals, a domain-specific discriminator is designed to align the feature distributions across the labeled source domain ($\mathbb{S}_{l}$), the unlabeled source domain ($\mathbb{S}_{u}$), and the target domain ($\mathbb{T}$). This discriminator aims to reduce the variability between these domains by ensuring that the extracted features from different sources are more consistent. By aligning these distributions, the domain-specific discriminator helps in minimizing the domain gap, thereby enhancing the model’s ability to generalize effectively to new and unseen data. The loss function of the domain-specific discriminator is given as
@@
Here, $\theta_{d}$ refers to the parameters of the domain classifier $d(\cdot)$. $p(X)$ is the domain label indicating whether $X$ is from the labeled source, unlabeled source, or target domain. $N$ denotes the number of samples in each batch. This loss function encourages the feature extractor to generate domain-invariant features by minimizing the ability of the domain classifier to distinguish between domains. As a result, the model learns to produce features that are robust across different domains, improving cross-subject generalization.
@@ IV-C5 Overall Loss Function
$\mathcal{L}=\alpha\mathcal{L}_{signal}+\beta\mathcal{L}_{gender}+\delta\mathcal{L}_{disc}+\mathcal{L}_{disease},$ (7)
where $\alpha$, $\beta$, and $\delta$ are weighting factors for the corresponding loss terms. The overall loss function ensures that the model learns robust, multi-dimensional EEG feature representations that are invariant to signal noise, sensitive to gender differences, adaptable across domains, and effective in classifying NSSI behaviors. By jointly optimizing these objectives, the model is better equipped to handle the complexities involved in EEG signal analysis for studies on depression and NSSI.
@@ V-A Implementation details and model setting
In the proposed NSSI-Net, the CNN encoder begins with an initial convolutional layer that employs 16 filters and a kernel size of 1 ×$\times\frac{\text{input\_size}}{2}+1$. This layer is followed by a depthwise convolution with 32 filters. Subsequent separable convolutions further refine these features, using kernel sizes of 1 ×$\times\frac{\text{input\_size}}{8}+1$. The BiGRU encoder is bidirectional, with each direction containing 128 units per layer, spread across 2 layers. Following the BiGRU layers, fully connected layers are employed to condense the feature representation, which systematically reduce and refine the feature dimensionality before passing it to the final classification and adversarial networks. The adversarial discriminators incorporate fully connected layers with ReLU activation functions.
@@ V-B Model Results under Semi-Supervised Learning
We compare the proposed NSSI-Net with various machine learning and deep learning models using a semi-supervised cross-subject ten-fold cross-validation protocol. As shown in Table I, the proposed NSSI-Net achieves superior performance, with an accuracy of 70.00±13.90. This represents a significant improvement of 7.44% compared to the best-performing model in the literature.
The table could be divided into two parts: traditional machine learning methods and deep learning methods. Among traditional machine learning approaches, Random Forest (RF [41]) achieves the highest accuracy at 55.39±07.04, followed closely by CORAL [42] with 57.07±03.45. Other traditional models, such as Support Vector Machines (SVM [43]) and Adaboost [44], demonstrate comparatively lower performance, with accuracies ranging from 48.09% to 57.85%. In contrast, deep learning methods offer more competitive results. The best performance in the literature is achieved by DDC [45], with an accuracy of 62.56±2.52. DAN [46] achieves an accuracy of 61.46±03.86, while DCORAL [47] reaches 60.91±02.95. These models outperform most traditional methods but still fall short compared to NSSI-Net. Another deep learning approach, DANN [48], records moderate performance with an accuracy of 58.20±02.80. Through a comprehensive comparison with the literature, the proposed NSSI-Net emerges as the top-performing model, surpassing both traditional and deep learning methods. Its 7.44% improvement over the next best model (DDC [45]) highlights its effectiveness in capturing the complex patterns associated with NSSI behaviors in adolescents, establishing it as a promising tool for early detection and intervention.
TABLE I: The mean accuracies (%) and standard deviations (%) of various traditional machine learning and deep learning methods for NSSI detection using semi-supervised cross-subject ten-fold cross-subject cross-validation. Models reproduced by the authors are indicated with an asterisk (*).
@@ V-C Analysis of Confusion Matrices
In summary, the confusion matrices reveal that while the model performs reasonably well overall, there are notable differences in performance across gender-specific subgroups. The model tends to be more accurate in identifying DN- cases in females and DN+ cases in males, which may be attributed to underlying biological and psychological differences between genders. However, the higher FP rate in males and the higher FN rate in females highlight areas for further improvement, particularly in addressing potential gender biases in the model’s predictions. Future improvements could focus on refining the gender-specific discriminator or incorporating a larger, more balanced dataset to enhance the model’s robustness and accuracy.

### v2

# NSSI-Net: A Multi-Concept GAN for Non-Suicidal Self-Injury Detection Using High-Dimensional EEG in a Semi-Supervised Framework

@@ Abstract
Non-suicidal self-injury (NSSI) is a serious threat to the physical and mental health of adolescents, significantly increasing the risk of suicide and attracting widespread public concern. Electroencephalography (EEG), as an objective tool for identifying brain disorders, holds great promise. However, extracting meaningful and reliable features from high-dimensional EEG data, especially by integrating spatiotemporal brain dynamics into informative representations, remains a major challenge. In this study, we introduce an advanced semi-supervised adversarial network, NSSI-Net, to effectively model EEG features related to NSSI. NSSI-Net consists of two key modules: a spatial-temporal feature extraction module and a multi-concept discriminator. In the spatial-temporal feature extraction module, an integrated 2D convolutional neural network (2D-CNN) and a bi-directional Gated Recurrent Unit (BiGRU) are used to capture both spatial and temporal dynamics in EEG data. In the multi-concept discriminator, signal, gender, domain, and disease levels are fully explored to extract meaningful EEG features, considering individual, demographic, disease variations across a diverse population. Based on self-collected NSSI data (n=114), the model’s effectiveness and reliability are demonstrated, with a 5.44% improvement in performance compared to existing machine learning and deep learning methods. This study advances the understanding and early diagnosis of NSSI in adolescents with depression, enabling timely intervention. The source code is available at https://github.com/Vesan-yws/NSSINet.
@@ II-B EEG Studies Exploring NSSI in Depressed Adolescents
In recent years, researchers have increasingly focused on utilizing EEG to explore the neural underpinnings of NSSI. In 2021, a study highlighted significant differences in EEG frequency and spatial components between depressive female adolescents who had attempted suicide and those who engaged in NSSI, demonstrating the potential of EEG as a diagnostic tool to distinguish between these two groups [21]. Moreover, subsequent research has revealed notable alterations in emotion-related EEG components and brain lateralization in response to negative emotional stimuli among adolescents with NSSI, suggesting that these neural changes could be crucial in understanding the emotional dysregulation associated with NSSI [22]. Further investigations have explored the effects of therapeutic interventions on EEG patterns in NSSI patients. For example, Zhao et al. ’s work observed significant changes in EEG microstate patterns before and after repetitive transcranial magnetic stimulation (rTMS) treatment, indicating that EEG can not only reflects the abnormal neural activity related to NSSI but also monitors the impact of therapeutic interventions [23]. These findings underscore the importance of EEG as a non-invasive, real-time measure that can provide insights into the neural dynamics associated with NSSI, potentially leading to more targeted and effective treatments.
@@ III-B Data Acquisition and Preprocessing
A standard EEG preprocessing pipeline was applied to the raw EEG signals to remove artificial artifacts, including eye movements, muscle activity, and environmental noise. Following artifact correction, the signals were downsampled from 500 Hz to 384 Hz. This reduction not only decreases data dimensionality but also retains essential temporal, spatial, and frequency information, thereby enhancing the efficiency of subsequent processing and analysis [32]. To increase the sample size for modeling, each 5-second trial was further segmented into multiple 1-second samples. As a result, the final data format can be represented as 114 participants $\times$ 35 trials $\times$ 5 seconds $\times$ 63 channels $\times$ 384 sampling points.
@@ IV-A Input Data
Considering the data imbalance (NSSI: 77 patients, including 65 females and 12 males; non-NSSI: 37 patients, including 18 females and 19 males) and the gender distribution, a data sampling strategy is implemented across $\mathbb{S}_{l}$, $\mathbb{S}_{u}$, and $\mathbb{T}$. As illustrated in Fig. 2, the sampling process for both the NSSI and non-NSSI groups is explained. For the NSSI group (denoted as DN+), 18 females are randomly selected from the 65 available candidates (red hatched), and all 12 males are included, ensuring equal gender representation within DN+. For the non-NSSI group (denoted as DN-), all 18 females (solid blue) are included, along with 12 randomly selected males from the 19 available candidates, achieving a balanced gender distribution comparable to the NSSI group. This sampling process results in a dataset of 30 NSSI patients (DN+) and 30 non-NSSI patients (DN-), balanced across both condition and gender.
@@ IV-B Spatial-Temporal Feature Extraction
To effectively extract spatial and temporal features from input EEG signals, we design a hybrid neural network that combines CNNs with GRUs. This architecture captures interactions both among different brain regions and across various time points, enhancing the representation of complex EEG dynamics.
@@
To further uncover the relationships of the brain patterns over time and encapsulate the key temporal dynamics of the entire EEG signal sequence, a BiGRU based encoder-decoder structure is designed. The BiGRU-based encoder for temporal information extraction is illustrated in Fig. 1, represented by the green vertical arrow. In our implementation, we use a BiGRU to process the sequence of feature vectors generated by the CNN encoder. The forward and backward passes of the BiGRU produce corresponding hidden states, which are concatenated to form the deep feature representation vector $\mathcal{A}=(a_{1},...,a_{M})\in\mathbb{R}^{M\times 1\times 2F}$. Once $\mathcal{A}$ is obtained from the concatenated hidden states of the BiGRU, it is fed into a Multi-Layer Perceptron (MLP) to derive $\mathcal{Y}=(y_{1},...,y_{M})\in\mathbb{R}^{M\times 1\times F}$. This MLP transforms the feature representation into a more refined feature space, capturing non-linear relationships between the elements of $\mathcal{A}$. Following this, a flatten operation is applied to $\mathcal{Y}$, resulting in the output vector $\mathcal{O}\in\mathbb{R}^{1\times MF}$. The output $\mathcal{O}$ is a rich representation that integrates both spatial features (extracted by the CNN layers) and temporal features (learned through the recurrent layers), providing a comprehensive understanding of the EEG signal’s dynamics.
@@ IV-C2 Gender-Specific Discriminator
Considering the gender bias present in NSSI patients and the variations in EEG representations between females and males, a gender-specific discriminator is designed to capture and leverage gender-specific EEG features. By integrating gender-specific discriminator, the model not only accounts for the gender-based differences but also improves its overall classification accuracy, ensuring that the subtleties in EEG signals attributed to gender are adequately represented and utilized for better prediction outcomes. The loss function of the gender-specific discriminator is given as
@@
Here, $\theta_{G}$ represents the parameters of the feature extractor (generator $G$). $\theta_{e}$ refers to the parameters of the gender classifier $e(\cdot)$. $\mathcal{O}$ refers to the features extracted as described in Section IV-B, where $\mathcal{O}^{\mathbb{M}}$ and $\mathcal{O}^{\mathbb{F}}$ represent the features of male and female samples, respectively. $N_{\mathbb{M}}$ and $N_{\mathbb{F}}$ denote the number of male and female samples in each batch, respectively. This loss function is designed to ensure that the model effectively captures gender-specific EEG features, which are crucial, particularly in addressing the gender bias present in the data.
@@ IV-C3 Domain-Specific Discriminator
Due to the inherent individual differences in EEG signals, a domain-specific discriminator is designed to align the feature distributions across the labeled source domain ($\mathbb{S}_{l}$), the unlabeled source domain ($\mathbb{S}_{u}$), and the target domain ($\mathbb{T}$). This discriminator aims to reduce the variability among these domains by ensuring that the extracted features from different sources are more consistent. By aligning these distributions, the domain-specific discriminator helps in minimizing the domain gap, thereby enhancing the model’s ability to generalize effectively to new and unseen data. The loss function of the domain-specific discriminator is given as
@@
Here, $\theta_{d}$ refers to the parameters of the domain classifier $d(\cdot)$. $p(X)$ is the domain label indicating whether $X$ is from the labeled source, unlabeled source, or target domain. $N$ denotes the number of samples in each batch. This loss function encourages the feature extractor to generate domain-invariant features, by minimizing the ability of the domain classifier to distinguish among domains. As a result, the model learns to produce features that are robust across different domains, improving cross-subject generalization.
@@ IV-C5 Overall Loss Function
${\mathcal{L}=\alpha\mathcal{L}_{signal}+\beta\mathcal{L}_{gender}+\delta\mathcal{L}_{disc}+\theta\mathcal{L}_{disease},}$ (7)
where $\alpha$, $\beta$, $\delta$ and $\theta$ are weighting factors for the corresponding loss terms. The overall loss function ensures that the model learns robust, multi-dimensional EEG feature representations that are invariant to signal noise, sensitive to gender differences, adaptable across domains, and effective in classifying NSSI behaviors. By jointly optimizing these objectives, the model is better equipped to handle the complexities involved in EEG signal analysis for studies on depression and NSSI.
@@ V-A Implementation details and model setting
A detailed summary of the model architecture, including the encoder and decoder configurations, is provided in Table I. The CNN encoder begins with an initial convolutional layer that employs 16 filters and a kernel size of 1 ×$\times\frac{\text{input\_size}}{2}+1$. This layer is followed by a depthwise convolution with 32 filters. Subsequent separable convolutions further refine these features, using kernel sizes of 1 ×$\times\frac{\text{input\_size}}{8}+1$. The BiGRU encoder is bidirectional, with each direction containing 128 units per layer, spread across 2 layers. Following the BiGRU layers, fully connected layers are employed to condense the feature representation, which systematically reduce and refine the feature dimensionality before passing it to the final classification and adversarial networks. The adversarial discriminators incorporate fully connected layers with ReLU activation functions.
@@
TABLE I: Summary of model architecture and configurations.
Layer (type:depth-idx) Kernel Shape Output Shape Param#
Input [384, 1, 63, 384]
Conv2d: 1-1 [1, 193] [384, 16, 63, 384] 3,104
BatchNorm2d: 1-2 [384, 16, 63, 384] 32
Conv2d: 1-3 [63, 1] [384, 32, 1, 384] 32,288
BatchNorm2d: 1-4 [384, 32, 1, 384] 64
MaxPool2d: 1-5 [1, 4] [384, 32, 1, 96]
Dropout: 1-6 [384, 32, 1, 96]
Conv2d: 1-7 [1, 49] [384, 32, 1, 96] 1,600
Conv2d: 1-8 [1, 1] [384, 16, 1, 96] 528
BatchNorm2d: 1-9 [384, 16, 1, 96] 32
MaxPool2d: 1-10 [1, 8] [384, 16, 1, 12]
Linear: 1-11 [384, 12, 16] 272
GRU: 1-12 [384, 12, 32] 3,264
Linear: 1-13 [384, 12, 16] 528
Linear: 1-14 [384, 12, 32] 544
GRU: 1-15 [384, 12, 32] 4,800
Linear: 1-16 [384,12,16] 528
MaxUnpool2d: 1-17 [1, 8] [384, 16, 1, 96]
ConvTranspose2d: 1-18 [1, 1] [384, 32, 1, 96] 544
ConvTranspose2d: 1-19 [1, 49] [384, 32, 1, 96] 1,600
BatchNorm2d: 1-20 [384, 32, 1, 96] 64
Dropout: 1-21 [384, 32, 1, 96]
MaxUnpool2d: 1-22 [1, 4] [384, 32, 1, 384]
ConvTranspose2d: 1-23 [63, 1] [384, 16, 63, 384] 32,272
BatchNorm2d: 1-24 [384, 16, 63, 384] 32
ConvTranspose2d: 1-25 [1, 193] [384, 1, 63, 384] 3,089
Total params: 164,807
Trainable params: 164,807
Non-trainable params: 0
Total mult-adds (G): 362.29
Input size (MB): 37.16
Forward/backward pass size (MB): 4958.66
Params size (MB): 0.34
Estimated Total Size (MB): 4996.16
@@ V-B Model Results under Semi-Supervised Learning
We compare the proposed NSSI-Net with various machine learning and deep learning models using a semi-supervised cross-subject ten-fold cross-validation protocol. In this framework, the labeled source domain $\mathbb{S}_{l}$ contributes to the optimization of all four discriminators: $\mathcal{L}_{signal}$, $\mathcal{L}_{gender}$, $\mathcal{L}_{domain}$ and $\mathcal{L}_{disease}$. The unlabeled source domain $\mathbb{S}_{u}$ and the target domain $\mathbb{T}$ contribute to the optimization of three discriminators: $\mathcal{L}_{signal}$, $\mathcal{L}_{gender}$, and $\mathcal{L}_{domain}$. Here, $\mathcal{L}_{disease}$ is trained in a supervised manner using the labeled data from $\mathbb{S}_{l}$. $\mathcal{L}_{signal}$, $\mathcal{L}_{gender}$, and $\mathcal{L}_{domain}$ leverage self-supervised learning to extract meaningful feature representations without requiring label information. As shown in Table II, the proposed NSSI-Net achieves superior performance, with an accuracy of 70.00±13.90. This represents a significant improvement of 5.44% compared to the best-performing model in the literature.
Specifically, the table could be divided into two parts: traditional machine learning methods and deep learning methods. Among traditional machine learning approaches, Adaboost [41] achieves the highest accuracy at 57.85±06.57, followed closely by CORAL [42] with 57.07±03.45. Other traditional models, including SA[43], RF[44], SVM [45], TCA [46], KPCA [47], and GFK [48], demonstrate comparatively lower performance, with accuracies ranging from 48.09% to 56.87%. On the other hand, the deep learning methods offer more competitive results. Among transfer learning approaches, DDC [49] achieves an accuracy of 62.56±02.52, followed by DAN [50] at 61.46±03.86, DCORAL [51] at 60.91±02.95, and DANN [52] at 58.20±02.80. Among existing methods, the best semi-supervised learning results are achieved by MixMatch [53] with an accuracy of 64.56$\pm$04.46, followed by PARSE [34] at 61.77$\pm$13.34. Through a comprehensive comparison with the literature, the proposed NSSI-Net proves to be the top-performing model, surpassing both traditional and deep learning methods. Its 5.44% accuracy improvement over the second-best model (MixMatch [53]) highlights its effectiveness in capturing the complex patterns associated with NSSI behaviors in adolescents, establishing it as a promising tool for early detection and intervention.
TABLE II: The mean accuracies (%) and standard deviations (%) of various traditional machine learning and deep learning methods for NSSI detection using semi-supervised cross-subject ten-fold cross-subject cross-validation. Models reproduced by the authors are indicated with an asterisk (*).
@@ V-C Analysis of Confusion Matrices
In summary, our current study reveals that the training of our gender-specific discriminator is likely constrained by two main factors: dataset imbalance and inherent biological and behavioral differences between genders. Due to the uneven gender distribution, the model may predominantly learn features from the female subgroup, which exhibits greater variance. Although we implement a data sampling strategy to minimize the impact of this imbalance, the lower variance in the male subgroup still leads to biased feature learning. Moreover, neurophysiological and behavioral markers associated with NSSI differ between genders. This variability makes it challenging for a uniform training process to fully capture the unique EEG patterns of each gender, thereby limiting the overall performance of the discriminator.